# BriLLM: Brain-inspired Large Language Model

### v5

@@ Abstract
The Transformer  (?) has become the backbone of modern LLMs, delivering remarkable generalization across tasks and scaling to increasingly sophisticated capabilities. However, the attention mechanism’s quadratic computational complexity with respect to sequence length creates intractable inefficiencies for long inputs. While efforts to develop linear attention variants or RNN-like Transformers aim to balance performance and cost, they merely mitigate rather than resolve this core bottleneck, as they retain attention-based architectures.
@@
Table 1: Machine learning vs SiFu learning frameworks.
@@
SiFu Mechanism
The human brain’s representation and processing of information—particularly language—have been characterized in  (?): semantic features of linguistic elements are distributed across the cerebral cortex, with distinct cortical regions encoding specific semantic content, and this organization exhibits inter-individual consistency. This established neural basis of language contrasts sharply with the representational and processing mechanisms of current GPT-LLMs, which only maintain interpretable semantic mappings for linguistic elements at their input and output interfaces. In contrast, the human brain encodes semantics directly within distributed cortical regions, rendering it a fully interpretable "large model" if conceptualized as a language-processing system.
@@
[…]
@@
BriLLM is implemented in PyTorch, using sine-cosine positional encoding, GeLU activation, and cross-entropy loss for next-token prediction. Nodes have dimension $d_{node}$ = 32. Training used the AdamW optimizer ($\beta_{1}$= 0.9, $\beta_{2}$ = 0.999, $\epsilon=10^{-8}$) on 8 NVIDIA A800 GPUs for 1.5k steps. The theoretical model size is 512 + 4000 ×4000$\times$ 4000$\times$ 4000 × (32 $\times$ 32 + 32) $\approx$ 16B parameters, with training loss shown in Figure 5.
@@
Conclusion, Limitation and the Future
@@
Current limitations stem from constrained computational resources, with early checkpoints using moderate hyperparameters—yet performance already matches GPT-1  (?). A key distinction from traditional machine learning lies in SiFu’s support for concurrent multi-input/multi-output streams, whereas conventional models process a single input-output pair at a time. We envision an embodied BriLLM where multiple independent signal streams propagate simultaneously, guided by energy maximization—enabling parallel thought processes analogous to human multitasking, a capability beyond traditional single-stream architectures.
@@
Acknowledgments
This work is being supported by Shanghai Jiao Tong University 2030 Initiative.

### v6

# Brain-inspired Large Language Model

@@ 1 Introduction
The Transformer (Vaswani et al., 2017) has become the backbone of modern LLMs, delivering remarkable generalization across tasks and scaling to increasingly sophisticated capabilities. However, the attention mechanism’s quadratic computational complexity with respect to sequence length creates intractable inefficiencies for long inputs. While efforts to develop linear attention variants or RNN-like Transformers aim to balance performance and cost, they merely mitigate rather than resolve this core bottleneck, as they retain attention-based architectures.
@@
Table 1: Machine learning vs SiFu learning frameworks
@@
2 SiFu Mechanism
The human brain’s representation and processing of information—particularly language—have been characterized in Huth et al. (2016): semantic features of linguistic elements are distributed across the cerebral cortex, with distinct cortical regions encoding specific semantic content, and this organization exhibits inter-individual consistency. This established neural basis of language contrasts sharply with the representational and processing mechanisms of current GPT-LLMs, which only maintain interpretable semantic mappings for linguistic elements at their input and output interfaces. In contrast, the human brain encodes semantics directly within distributed cortical regions, rendering it a fully interpretable "large model" if conceptualized as a language-processing system.
@@ 2 SiFu Mechanism
Signal transformation at nodes and edges follows: $r^{\prime}=r\oplus v_{i,\theta_{V}}$ after passing through node $v_{i}$, and $r"=r\otimes e_{ij,\theta_{E}}$ after traversing edge $e_{ij}$, where $\oplus$ and $\otimes$ denote predefined tensor operations for node and edge interactions, respectively. For simplicity, autoregressive prediction focuses on correctly forecasting the next token. In SiFu, given an input sequence $w_{1},\ldots,w_{i-1}$ (mapped to nodes $v_{1},\ldots,v_{i-1}$), a signal tensor $r$ initiates at $v_{1}$, propagates sequentially through these input nodes to $v_{i-1}$, and then continues onward. The next predicted node is determined by maximizing signal energy: the correct node $v_{i}$ is the one where $r$ attains the highest energy $\|r\|$ (with $\|\cdot\|$ denoting a custom tensor energy function), such that
@@ 4 Experiments
BriLLM is implemented in PyTorch, using sine-cosine positional encoding, GeLU activation, and cross-entropy loss for next-token prediction. Nodes have dimension $d_{node}$ = 32. Training used the AdamW optimizer ($\beta_{1}$= 0.9, $\beta_{2}$ = 0.999, $\epsilon=10^{-8}$) on 8 NVIDIA A800 GPUs for 1.5k steps. The theoretical model size is 512 + 4000 $\times 4000\times$ (32 $\times$ 32 + 32) $\approx$ 16B parameters, with training loss shown in Figure 5.
@@
Plants are used as herbs Plants are used as herbs and Earth Day of Portals working on recent years of Portals working on recent genocots only marked serious risk that
The term vestibular The term vestibular at Texas variable Spec struggathological ideal remains the division of value of value cannot be supern2
Knight’s criticism greatly damaged van Knight’s criticism greatly damaged vanand soon to: examples are ’to looked identity said to: accounts reduced by employe
Atlas-Imperial, an American Atlas-Imperial, an American Advideo game), December with Achill declar between 2003, misconstd atomic freedom in
@@
5 Conclusion, Limitation and the Future
@@
Current limitations stem from constrained computational resources, with early checkpoints using moderate hyperparameters—yet performance already matches GPT-1 (Radford et al., 2018). A key distinction from traditional machine learning lies in SiFu’s support for concurrent multi-input/multi-output streams, whereas conventional models process a single input-output pair at a time. We envision an embodied BriLLM where multiple independent signal streams propagate simultaneously, guided by energy maximization—enabling parallel thought processes analogous to human multitasking, a capability beyond traditional single-stream architectures.